OpenStudy (anonymous):

Hayden invested $750 for three years at an interest rate of 0.035. How much interest did he earn in

OpenStudy (wolf1728):

The formula is Total = principal * (1+rate)^years Total = 750 * (1.035)^3 Total = 750*1.108717875 Total = 831.54